OpenStudy (anonymous):

Okay, I'm having some issues understanding my last homework assignment for the week... Tickets=-0.2^2+12x+11 (x=1 is the day tickets go on sale) does the graph of this equation open up or down?

OpenStudy (anonymous):

oh my there should be a template made for this problem because there is at least one a day. where does it come from?!

OpenStudy (anonymous):

it is a parabola that faces down because the "leading coefficient" is -.2 and that is negative

OpenStudy (anonymous):

so it "opens down"

OpenStudy (anonymous):

are you missing an X? should it be -0.2x^2?

OpenStudy (anonymous):

to find the maximum number of tickets sold, and the day on which they are sold, use \[-\frac{b}{2a}\] in this case \[a=-.2,b=12\] so \[-\frac{b}{2a}=\frac{12}{.4}=\frac{120}{4}=30\]

OpenStudy (anonymous):

yeah I was missing an 'x' sorry I have sticky keys

OpenStudy (anonymous):

it should open downward then :)

OpenStudy (anonymous):

maximum tickets are sold on day 30. to find out how many tickets are sold, replace x by 30 and get \[-.2\times 30^2+12\times 30+11=191\]

OpenStudy (anonymous):

so on day 30 you sell 191 tickets and that is the maximum. it is also the vertex of the parabola. the vertex is (30,191)

OpenStudy (anonymous):

there may be another question. like solve \[-0.2x^2+12x+11 =0\] i forget. last day of ticket sales maybe?

OpenStudy (anonymous):

after how many days will the peak/low occur... lol